OpenStudy (anonymous):

I need help factoring w^2-8w=0.

OpenStudy (anonymous):

w*(w-8) = 0

OpenStudy (anonymous):

w(w-8)=0 w=0 w-8=0 w=8 answer is either 0 or 8

OpenStudy (anonymous):

Thank y'all!